Nicholas21215:

A marketing team is going to buy promotional coffee mugs as giveaways. The manufacturing company charges a fixed $250 setup fee and $8 per mug.

Eveeeeeeeee14:

250-8=242

Eveeeeeeeee14:

there you go

Cyber:

@eveeeeeeeee14 wrote:
250-8=242
You are not allowed to give out direct answers. But you can try to help them come up with an answer on their own.

Cyber:

Just giving you an FYI for the future.

Cyber:

This post can be closed now.

kacchan:

@cyber wrote:
This post can be closed now.
There offline so they can't close the post!

Cyber:

@kacchan wrote:
@cyber wrote:
This post can be closed now.
There offline so they can't close the post!
Any Mod Online Can close this

kacchan:

@cyber wrote:
@kacchan wrote:
@cyber wrote:
This post can be closed now.
There offline so they can't close the post!
Any Mod Online Can close this
i know

Batman567:

@eveeeeeeeee14 wrote:
250-8=242
you have to help theme get the answer don't give him or her the answer that called direct answer

TETSXPREME:

Welcome To Questioncove! What is the question that you are figure out ?